Doug is 12 years older than his sister Linda. in 4 years ,he will be twice as old as linda. how old is each of them now?
Answer by edjones(8007)

Let d=Doug's age, n=Linda's age
d=n+12...d-n=12 (A)
d+4=2(n+4)...d+4=2n+8...d-2n=4 (B)
.
d-n+12
-d+2n=-4 multiply (B) yimes -1
--------
n=8 add (A) and (B)
Linda is 8 and Doug is 20
in 4 yrs Linda will be 12 and Doug 24.
.
Ed